OpenStudy (k.hashimi):

There are two solutions to the equation \log(30 x) + \log(x-1) = 1log(30x)+log(x−1)=1. Find the larger of the two solutions.

jhonyy9 (jhonyy9):

log 30x + log (x-1) = 1 there are two samw equations so why ?

jhonyy9 (jhonyy9):

same - equations - sorry

jhonyy9 (jhonyy9):

do you know this property of logarithms ? log a + log b = log a*b

jhonyy9 (jhonyy9):

can using this in case of your exercise ?

OpenStudy (k.hashimi):

it just got repeated sorry.

jhonyy9 (jhonyy9):

ok np but can using this formula in case of your exercise ?

jhonyy9 (jhonyy9):

a=30x and b=(x-1)

OpenStudy (k.hashimi):

u can use anything to solve it but yeah iam familiar with it

OpenStudy (k.hashimi):

but what im confused abut is the 2 solutions that they are reffering to

jhonyy9 (jhonyy9):

using this formula you get log 30x +log (x-1) =1 log [30x(x-1)] = 1 log (30x^2 -30x) =1 do you know how many is the base of this logarithm noted by log ?

jhonyy9 (jhonyy9):

Loser how many is the base of this log. just bc. i ve learned it 2 by 35 years ago here in Europe and here in OS have wrote me that is 10 ?

jhonyy9 (jhonyy9):

ok- ty

jhonyy9 (jhonyy9):

so in this case we can rewriting the 1 like log 10

jhonyy9 (jhonyy9):

and will get this equation log (30x^2 -30x) =1 log (30x^2 -30x) = log 10 30x^2 -30x = 10 do you can continue it now ?

jhonyy9 (jhonyy9):

subtract from both sides 10 and will get 30x^2 -30x -10 = 0 can you solving this quadratic ?

jhonyy9 (jhonyy9):

factorize out the 10 and will get 10(3x^2 -3x -1) = 0 divide both sides by 10 and will get 3x^2 -3x -1 = 0 can you solving it now for x ?

jhonyy9 (jhonyy9):

do you know formula of discriminant ?

jhonyy9 (jhonyy9):

please collaborate

OpenStudy (k.hashimi):

i was trying to solve the last equation you wrote. sorry

OpenStudy (k.hashimi):

yeah i do

jhonyy9 (jhonyy9):

3x^2 -3x -1 = 0 so can you solving this quadratic ?

OpenStudy (k.hashimi):

let me solve it and confirm wiht u

jhonyy9 (jhonyy9):

ok. but please write here your work

jhonyy9 (jhonyy9):

discriminant D = ?

jhonyy9 (jhonyy9):

step by step please

OpenStudy (k.hashimi):

ok i just solved it using the formula b+/= ....

OpenStudy (k.hashimi):

and i got 2 values for x x=1.26 x= -0.263

OpenStudy (k.hashimi):

i dont know if thats right or wrong

jhonyy9 (jhonyy9):

D = b^2 -4ac is the formula of discriminant

jhonyy9 (jhonyy9):

3x^2 -3x -1 = 0 how many you have got the value for discriminant ?

OpenStudy (k.hashimi):

21

jhonyy9 (jhonyy9):

D = 9 +12 = 21 yes is right so than x_1,2 = ?

OpenStudy (k.hashimi):

im confused what are using d for

jhonyy9 (jhonyy9):

D is the discriminant

jhonyy9 (jhonyy9):

after you get the value of discriminant how you get the roots x_1,2 ?

OpenStudy (k.hashimi):

idk

jhonyy9 (jhonyy9):

so how you ve got these two values ?

jhonyy9 (jhonyy9):

ok. try using this x_1,2 = (-b +/- sqrt(D) )/2a

jhonyy9 (jhonyy9):

3x^2 -3x -1 = 0 b = ?

jhonyy9 (jhonyy9):

D = b^2 -4ac = ?

jhonyy9 (jhonyy9):

how you get the value of discriminant ?

jhonyy9 (jhonyy9):

3x^2 -3x -1 = 0 a=3 b=-3 c=-1 yes ?

jhonyy9 (jhonyy9):

bc. there are ax^2 +bx +c = 0

jhonyy9 (jhonyy9):

so can you calculi now the x_1,2 = ?

jhonyy9 (jhonyy9):

x_1,2 = (-b +/- sqrt(D) )/2a

OpenStudy (k.hashimi):

ok ur method is different